OpenStudy (anonymous):

The normal to the curve y = x^2+3x+1 at the point (-1,1) cuts the curve again at Q. Find the coordinates of Q. I think the question is flawed, or I am really retarded...

ganeshie8 (ganeshie8):

find the normal first

OpenStudy (anonymous):

Its 1 right?

ganeshie8 (ganeshie8):

you have a typo in the quesiton i think. the point should be (-1, -1) i think

OpenStudy (anonymous):

After m1m2 = -1, where m1=-1

ganeshie8 (ganeshie8):

the given point wont even exist on the given curve :o

OpenStudy (anonymous):

I knew it! This is a question my teacher gave, I knew it was wrong

ganeshie8 (ganeshie8):

good :) may be solve it for (-1, -1) and tell her the typo...

OpenStudy (anonymous):

Thankyou so much sir! Much appreciated!

ganeshie8 (ganeshie8):

np :) yes the slope of tangent at (-1, -1) is 1 so the normal of slope wud be -1

ganeshie8 (ganeshie8):

*slope of normal wud be -1

OpenStudy (anonymous):

Yep thats true :)

ganeshie8 (ganeshie8):

write out the equation of normal and solve it wid the curve

OpenStudy (anonymous):

Thankyou so much!

ganeshie8 (ganeshie8):

yw :)

OpenStudy (anonymous):

Hey Ganesha, once we get the normal how do we solve it with the curve?

ganeshie8 (ganeshie8):

you got the normal curve :- line thru point (-1, -1) wid a slope of -1 y+1 = -1(x+1) ------------------------------(1) right ?

OpenStudy (anonymous):

Yes, and I eventually got 0 = - x - y - 2 What do i do next?

ganeshie8 (ganeshie8):

your equation is, y = x^2+3x+1 -------------------(2)

ganeshie8 (ganeshie8):

solve (1) and (2) that gives u the intersection points of normal and given curve

OpenStudy (anonymous):

But for (2) do i use the dy/dx of it or just that equation?

ganeshie8 (ganeshie8):

just the equation, cuz you want to find intersection between curve and normal

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

Ok, but how do I find the intersection when one equation has power and one doesnt?

ganeshie8 (ganeshie8):